Determine the concentration of a solution prepared by diluting 20.0 mL of 2.00 M NaCl to 250.0 mL
Answer #1

using the molarity equation,

M1V1=M2V2

given, initial concentration(M1): 2.00 M

initial volume: 20.0 ml

final volume(V2): 250.0 ml

final concentration: let M2

2*20 = M2*250

M2= 40/250

M2= 0.16 M

so, final concentration after dilution is 0.16M.